# 11footnotetext: Corresponding authors22footnotetext: Project lead

### v1

@@ 5.3 Ablation Studies
In this section, we study the effectiveness of different components in the proposed IntLoRA. The experiments are conducted with IntLoRASHIFT on the subject-driven generation with 15 subject-prompt pairs under 4-bit weight quantization.
Contradictory aspect of auxiliary variances. In Sec. 4.3, we have theoretically pointed out that there is a choice dilemma for $\sigma_{\mathbf{R}}$. Here we elaborate on its effect through distribution visualization. Specifically, we remove the VMC and use a scaling scalar to generate a too-large or too-small auxiliary variance, followed by the log2 quantization on the adaptation term. The results are shown in Fig. 6. On the one hand, setting $\sigma_{\mathbf{R}}$ too large can lead to a low correlation $\rho(\mathbf{W},\mathbf{W-R})$, which makes it hard to reconstruct $\mathbf{W}$ from $\mathbf{W-R}$ using estimator $\mathbf{W}\approx\mathcal{Q}(\mathbf{W-R})+\mathbf{R}$. On the other hand, a too small $\sigma_{\mathbf{R}}$ prevents the expectation of adaptation term converging to zero, causing few log bins to be used. In experiments, we find that the training of both settings fails to converge. By contrast, the proposed VMC can precisely control $\sigma_{\mathbf{R}}$ to allow most values of the adaptation term to be zero-neighbored, facilitating more challenging log2 quantization. Moreover, it should be noted that the too small $\sigma_{\mathbf{R}}$ can also be regarded as an approximation of direct quantization on the zero-initialized $\mathbf{AB}$, and thus the experimental results also justify the AQS for zero-initialized $\mathbf{AB}$.
Table 5: Ablation on different distribution choices of the auxiliary matrix.
@@
StudentT 0.2935 0.6420 0.2587 0.8048
Laplace 0.3992 0.6680 0.2548 0.8311
Distribution selection for auxiliary matrix. In the proposed IntLoRA, the auxiliary matrix $\mathbf{R}$ plays a crucial role in both AQS and VMC. Therefore, the distribution shape of $\mathbf{R}$ can greatly influence the performance. To this end, we ablate different distributions as the instantiation of $\mathbf{R}$ and give the corresponding performances in  Tab. 5. It can be seen that the Laplace distribution performs better than the other options. This is because a light-tailed distribution, such as Laplace, clusters most samples around zero, which is consistent with the requirement for log2-quantization. And this accurate log2-quantization further facilitates favorable performance. Therefore, the light-tailed distribution performs experimentally better than the heavy-tailed counterparts. For better effects, we provide the shape of different distributions in  Fig. 12.
@@
Figure 7: The normalized performance under different $\alpha$.
Ablation on the smoothing factor. In order to produce quantization-friendly adaptation term, we propose the VMC which rescales $\sigma_{\mathbf{R}}$ by $r^{\alpha}$, where $r$ is the variance ratio $\frac{\sigma_{\mathbf{W}}}{\sigma_{\mathbf{R}}}$ and $\alpha$ is a hyperparameter to search for the optimal control. Given $r<1$ (the distribution of $r$ is given in Fig. 10), making $\alpha<1$ further scales up $\sigma_{\mathbf{R}}$, which leads to lowered correlation between $\mathbf{W}$ and $\mathbf{W-R}$, making it difficult to recover $\mathbf{W}$ from $\mathbf{W-R}$. Moreover, setting $\alpha>1$ can decrease $\sigma_{\mathbf{R}}$, making the distribution being biased according to Eq. 5. Experiments in Fig. 7 show that setting $\sigma_{\mathbf{R}}$ slightly smaller than $\sigma_{\mathbf{W}}$ can obtain favorable performance, indicating that the information loss of $\mathbf{W}$ has a greater impact than a biased adaptation term. In the implementation, we chose a moderate smoothing factor $\alpha=1.5$ for the trade-off.
6 Conclusion
We propose IntLoRA, a novel joint adaptation and quantization framework which can address the arithmetic inconsistency by employing integer low-rank parameters, to push the efficiency limits of diffusion model fine-tuning. We propose the quantization-adaptation separation to allow the coexistence of zero-initialized gradient and quantization-friendly distribution. We further introduce the multiplicative low-rank adaptation to achieve a decoupled quantizer of pre-trained and adaptation weights for PTQ-free weight merge, accompanied by the variance matching control to adjust the channel-aware variance for accurate adaptation control. Benefiting from the elegant design, we provide two variants of IntLoRA, which either use int-multiplication or bit-shifting to adapt the quantized pre-trained models. Through transferring the adaptation weights to the integer arithmetic, our IntLoRA demonstrates its effectiveness across different pre-trained models and various downstream tasks, while exhibiting impressive efficiency across model tuning, storage, and inference.
Appendix
Appendix A Evaluation on natural language
In the main paper, we mainly validate the effectiveness of our IntLoRA on the image generation tasks. In this section, we turn our focus to the natural language to give a broader view of performance comparison. Specifically, we apply different methods to fine-tune the Llama3-8B model (Dubey et al., 2024) and use the MetaMathQA (Yu et al., 2023) dataset for training and GSM8K (Cobbe et al., 2021) dataset for testing. In all experiments, we fine-tune the model using the AdamW optimizer for 2 epochs with the cosine learning scheduler, and the warm-up ratio of the learning scheduler is set as 0.005. We follow the evaluation tools in MetaMathQA, which use the Alpaca prompt and evaluate the model in zero-shot. Due to the GPU memory limitation, we only append the LoRA layer to the attention query and value projection matrices. The comparison results are shown in Sec. H.1. It can be seen that our IntLoRA matins stable performance when transferring to natural language tasks. We also provide some testing cases in Appendix H for better presentation.
Appendix B Further explanation of the auxiliary matrix.
In the proposed method, the auxiliary matrix $\mathbf{R}$ is crucial to ensure quantization-friendly distribution and meaningful optimization trajectories. Here, we summarize two aspects of the role of $\mathbf{R}$ in the proposed method. First, $\mathbf{R}$ can be viewed as an initialization parameter for trainable parameters $\mathbf{AB}$. Specifically, we use $\mathbf{R}$ without gradient in the proposed AQS to guarantee easily quantized adaptation distributions while guaranteeing zero-initialized gradient trajectories. Second, the presence of $\mathbf{R}$ in the adaptation term also implies that it can be used to control the distribution of the adaptation term, thus allowing for more challenging quantizers, such as log2 quantization. For this, we propose the VMC, which makes the expectation of the adaptation term converge to zero, enabling as many log-scale buckets to be used as possible. Despite this indispensable role of $\mathbf{R}$, finding the optimal shape and variance of $\mathbf{R}$ remains challenging. In this work, we attempt to solve this from an engineering perspective by ablating different distributions and the variance scales. Experimental results demonstrate the effectiveness of this solution.
Appendix C Differences from QA-LoRA
In this section, we give a detailed discussion about the differences between the proposed IntLoRA and QALora (Xu et al., 2023b). First, one of the limitations of QA-LoRA is that it requires the pre-trained weights and the adaptation weights to share the same quantizer, which leads to a discrete parameter search space, limiting the adaptation capability for downstream tasks. In contrast, our IntLoRA ensures a more powerful adaptation through the proposed MLA which reformulates the form from additive to multiplicative, thus allowing independent quantizers. Second, the adaptation weights of QALoRA are still in FP, leading to storage cost as well as FP addition for weight merge during deployment, whereas our IntLoRA only needs to store INT weights while fusing weights through hardware-efficient integer multiplication or bit shifting. Third, QALoRA requires additional group quantization in the input channel dimensions, leading to additional GEMM core design, while our approach uses a common quantization strategy, thus facilitating practical deployment.
Appendix D More implementation details
For the subject-driven generation, we use the AdamW optimizer with a weight decay of 1e-2 and fine-tune the query, key, value, and output projection layer. The learning rate is set to 6e-5. The batch size is set to 1, and the number of training steps is 400. The rank of the LoRA is set to 4. We adopt the prior preservation strategy as Dreambooth (Ruiz et al., 2023) to generate 200 class images. For the controllable generation, we fine-tune the model for 11 epochs for Canny-to-Image tasks and 20 epochs for Landmark-to-Face and Segmentation-to-Image tasks. The learning rate is set to 1e-5 using the AdamW optimizer. The LoRA rank is set to 4. The batch size is set to 8 and the image resolution is $512\times 512$ for all three tasks. For the style-customized generation, we fine-tune the pre-trained model using the AdamW optimizer with a learning rate of 5e-5. Since it involves a larger SDXL, we chose a relatively large LoRA rank of 64 for all compared methods, since there is only one style image as well as the larger pre-trained parameters. We fine-tune for 500 steps with batch size 1. Similar to StyleDrop (Sohn et al., 2023), we only use one image as the style image and find it works well. The style images and text prompts for evaluation are given in Appendix H. The variance ratio in the variance matching control is surrogated as the value range ratio, i.e., $r=\frac{\max\{|\max(\mathbf{W})|,|\min(\mathbf{W})|\}}{\min\{|\max(\mathbf{R})|,|\min(\mathbf{R})|\}}$.
Table 6: Quantitative comparison on the natural language tasks of mathematical answering.
Methods LoRA-FP QLoRA QA-LoRA IR-QLoRA IntLoRAMUL IntLoRASHIFT
#praram 1.17M 1.17M 1.18M 1.18M 1.17M 1.17M
accuracy 62.24% 64.06% 63.53% 63.00% 64.23% 64.10%
Appendix E Additional ablation experiments
@@
Ablation on the LoRA rank. The low-rank $d$ in LoRA is a trade-off between performance and efficiency. A larger rank improves the adaptation ability by training more parameters but comes with larger training and storage costs, and vice versa. Here, we give the impact of different rank setups on performance in  Fig. 8. One can see that the performance generally improves as we increase the rank, but the rate of growth varies. For instance, the increase speed from rank=4 to rank=8 increases inferior to the one of from rank=2 to rank=4. Moreover, increasing the rank to 16 can generally obtain better results than its lower counterpart. In practice, considering the trade-off between performance and efficiency, we select a moderate rate rank=4.
@@
The effects of variance matching control for IntLoRAMUL. In this work, we propose the variance matching control to adjust the variance of $\mathbf{R}$, so that allows the log2-quantization of the adaptation term to obtain the IntLoRASHIFT. In other words, the VMC is primarily introduced for IntLoRASHIFT. Despite we also apply the VMC to IntLoRAMUL, given the IntLoRAMUL does not require such strict constraints on the distribution shape of the adaptation term, it is interesting to investigate the influence of variance matching control on the performance of IntLoRAMUL. To this end, we adjust the smoothing factor $\alpha$ to adjust the strength of the VMC, e.g., setting $\alpha$ to zero can lead to the removal of the VMC. The results of the IntLoRAMUL under different VMC scales are shown in  Fig. 9. As one can see, despite the VMC being initially proposed for the log2-quantization, the well-structured distribution also facilitates uniform quantization. For example, when we set the $\alpha$ approaching to zero, i.e. the VMC is close to being removed, the performance of IntLoRAMUL appears similar pattern as the IntLoRASHIFT, which suffers a significant performance drop. Moreover, the performance gains gradually converge when the $\alpha>1.5$. In short, the VMC can not only allow the log2-quantization to work but also improve the performance of the uniform quantization.

### v3

@@ 5.4 Ablation Studies
Ablation on the smoothing factor. As discussed in Sec. 4.1, there is a dilemma in choosing an appropriate $\sigma_{\mathbf{R}}$. For example, setting it too large can lead to information loss of the original weights, while a too-small one results in a large quantization error. To this end, we introduce $r^{\alpha}$ in the proposed VMC as the hyperparameter to search for the task-oriented variance. We give the downstream task performance with varying $\alpha$ in Fig. 7. It can be seen that setting $\sigma_{\mathbf{R}}$ slightly smaller than $\sigma_{\mathbf{W}}$ can obtain better performance, indicating that the information loss has a greater impact than the quantization error. In the implementation, we chose a moderate smoothing factor $\alpha=1.5$ for the trade-off.
Distribution selection for auxiliary matrix. In this work, the auxiliary matrix $\mathbf{R}$ plays a crucial role in both AQS and VMC. Therefore, the distribution shape of $\mathbf{R}$ can potentially influence the performance. To this end, we investigate different distribution shapes of $\mathbf{R}$ through ablation experiments and give the results in  Tab. 4. It can be seen that the Laplace distribution performs better than other options on most metrics. This is because a light-tailed distribution, such as Laplace, clusters most samples around zero, which facilitates smaller errors for log2-quantization. Therefore, the light-tailed distribution performs empirically better than its heavy-tailed counterparts.
Figure 6: Qualitative comparison on style-customized generation. The text prompt is “A friendly robot in [V] style”, “A panda eating bamboo in [V] style”, and “The letter ‘G’ in [V] style”, respectively. “Ours” denotes the IntLoRASHIFT. More results are provided in Appendix H.
Table 4: Ablation on different distribution shape choices of the auxiliary matrix.
@@
StudentT 0.2935 0.6420 0.2487 0.8048
Laplace 0.4492 0.6980 0.2588 0.8110
Figure 7: The normalized performance under different $\alpha$.
6 Further Discussion
Results on NLP tasks. In addition to fine-tuning diffusion models for image generation, we further validate the effectiveness of the NLP tasks. Specifically, we fine-tune the Llama3-8B model (Dubey et al., 2024) and use the MetaMathQA dataset (Yu et al., 2023) for training and GSM8K dataset (Cobbe et al., 2021) for testing. The comparison results are shown in Tab. 5. It can be seen that our IntLoRA maintains stable performance when transferring to natural language. For example, IntLoRAMUL outperforms QLoRA by 0.17% QA accuracy under 8-bit quantization, demonstrating the generalization of our IntLoRA.
Difference from EfficientDM. EfficientDM (He et al., 2023) employs QAT-like LoRA fine-tuning on the FP16 diffusion weights for network quantization. Despite it is inference-efficient as it can directly produce quantized merged weights, we would like to point out that it is not training-efficient. Specifically, EfficientDM requires load FP16 pre-trained weights at the training stage, which is unacceptable for fine-tuning large-size models on consumer-level GPUs. By contrast, our IntLoRA is both training and inference efficient. Moreover, we also explore our IntLoRA on the diffusion quantization task. The results are shown in Tab. 6. It can be seen that our IntLoRAMUL achieves even better performance than the EfficientDM. It should be noted that our IntLoRA only needs to load the quantized weights during calibration instead of the floating-point weights in EfficientDM, thus reducing the training memory cost.
Table 5: Comparison on the natural language task of mathematical answering. More qualitative results in Appendix H.
Methods LoRA QLoRA QA-LoRA OursSHIFT OursMUL
nbits W16A16 W8A8 W8A8 W8A8 W8A8
accuracy 64.24% 64.06% 63.53% 64.10% 64.23%
Table 6: Comparison with EfficientDM on W4A4 diffusion model quantization. We evaluate on the ImageNet $256\times 256$ image generation, and train with ddim_step=20 on LDM-4 model with 500 training epochs.
[…]
EfficientDM 178.20 13.42 26.67 0.70
OursSHIFT 116.50 20.20 26.79 0.63
OursMUL 199.20 10.43 24.02 0.79
7 Conclusion
We propose IntLoRA, which employs integer-type low-rank parameters, to remove the additional PTQ on the merged weights. Specifically, we introduce the quantization-adaptation separation to allow the coexistence of zero-initialized gradient and quantization-friendly distribution. We further develop the multiplicative low-rank adaptation to achieve a decoupled quantizer of pre-trained and adaptation weights, accompanied by the variance matching control to adjust the variance for accurate adaptation control. Benefiting from these elegant designs, we provide two variants of IntLoRA, which either use int-multiplication or bit-shifting to adapt the quantized pre-trained models. Through transferring the adaptation weights to the integer arithmetic, our IntLoRA demonstrates its effectiveness across different pre-trained models and various downstream tasks, while exhibiting impressive both training and inference efficiency.
Acknowledges
This work is supported in part by the National Natural Science Foundation of China, under Grant (62302309, 62171248), Shenzhen Science and Technology Program (JCYJ20220818101014030, JCYJ20220818101012025).
Impact Statement
This work aims to achieve efficient fine-tuning of quantized diffusion models, reducing both training and inference costs without sacrificing performance. It has no ethical concerns and can lower the resource barrier for model customization, thus enabling broader access to generative AI.
Appendix A Summery of IntLoRA Algorithm
Before tuning, we pre-process the pre-trained weights in Algo. 1, followed by the forward process of IntLoRAMUL and IntLoRASHIFT in Algo. 2 and Algo. 3, respectively.
Algorithm 1 The weight pre-process of the linear layer in IntLoRA
  Input: Pre-trained wight $\mathbf{W}\in\mathbb{R}^{C_{out}\times C_{in}}$, auxiliary matrix $\mathbf{R}\in\mathbb{R}^{C_{out}\times C_{in}}$, smooth factor $\alpha\in\mathbb{R}$
  Output: Quantitzed pre-trained weights $\mathbf{W}_{\mathrm{round}}$, scaling factor $s_{\mathrm{round}}$, zero point $z_{\mathrm{round}}$
  sigma_R $\leftarrow$ variance estimation of $\mathbf{R}$
  sigma_W $\leftarrow$ variance estimation of $\mathbf{W}$
  r = (sigma_W / sigma_R)α
  $\mathbf{R}_{\mathrm{star}}$ = r * $\mathbf{R}$
  $\mathbf{W}_{\mathrm{process}}=\mathbf{W-R_{\mathrm{star}}}$
  $\mathbf{W}_{\mathrm{round}},s_{\mathrm{round}},z_{\mathrm{round}}\leftarrow$ uniform_quantizer($\mathbf{W}_{\mathrm{process}}$)
Algorithm 2 The forward process of the linear layer in IntLoRAMUL
  Input: Pre-processed quantized weights $\mathbf{W}_{\mathrm{round}}$, scaling factor $s_{\mathrm{round}}$, zero point $z_{\mathrm{round}}$, auxiliary matrix $\mathbf{R}_{\mathrm{star}}\in\mathbb{R}^{C_{out}\times C_{in}}$, LoRA parameters $\mathbf{A}\in\mathbb{R}^{C_{out}\times d}$, $\mathbf{B}\in\mathbb{R}^{C_{out}\times d}$, input tensor $\mathbf{x}\in\mathbb{R}^{C_{in}\times L}$
  Output: Output tensor $\mathbf{y}\in\mathbb{R}^{C_{out}\times L}$
  $\mathbf{W}_{\mathrm{adapt}}=s_{\mathrm{round}}\cdot\mathbf{I}+\frac{1}{\mathbf{W}_{\mathrm{round}}-z_{\mathrm{round}}}\odot(\mathbf{R_{\mathrm{star}}+AB})$
  $\mathbf{W}_{\mathrm{adapt}}^{\texttt{INT}},s_{\mathrm{adapt}},z_{\mathrm{adapt}}\leftarrow\text{uniform\_quantizer}(\mathbf{W}_{\mathrm{adapt}})$
  $\mathbf{x}^{\texttt{INT}},s_{\mathrm{x}}\leftarrow\text{act\_quantizer}(\mathbf{x})$
  $\mathbf{W}_{\mathrm{merge}}=(\mathbf{W}_{\mathrm{adapt}}^{\texttt{INT}}-z_{\mathrm{adapt}})\odot(\mathbf{W}_{\mathrm{round}}-z_{\mathrm{round}})$
  $\mathbf{y}=s_{\mathrm{x}}s_{\mathrm{round}}\mathbf{W}_{\mathrm{merge}}\mathbf{x}^{\texttt{INT}}$
Algorithm 3 The forward process of the linear layer in IntLoRASHIFT
  Input: Pre-processed quantized weights $\mathbf{W}_{\mathrm{round}}$, scaling factor $s_{\mathrm{round}}$, zero point $z_{\mathrm{round}}$, auxiliary matrix $\mathbf{R}_{\mathrm{star}}\in\mathbb{R}^{C_{out}\times C_{in}}$, LoRA parameters $\mathbf{A}\in\mathbb{R}^{C_{out}\times d}$, $\mathbf{B}\in\mathbb{R}^{C_{out}\times d}$, input tensor $\mathbf{x}\in\mathbb{R}^{C_{in}\times L}$, desired bit-width $b$, pre-defined max bit-width number $N=32$
  Output: Output tensor $\mathbf{y}\in\mathbb{R}^{C_{out}\times L}$
  $\mathbf{W}_{\mathrm{adapt}}=s_{\mathrm{round}}\cdot\mathbf{I}+\frac{1}{\mathbf{W}_{\mathrm{round}}-z_{\mathrm{round}}}\odot(\mathbf{R_{\mathrm{star}}+AB})$
  shift = $\text{clip}(\lfloor-\log_{2}|\mathbf{W}_{\mathrm{adapt}}|\rceil,0,2^{b}-1)$
  $\mathbf{W}_{\mathrm{merge}}=\frac{1}{2^{N}}\odot\operatorname{sign}(\mathbf{W}_{\mathrm{adapt}})\odot[(\mathbf{W}_{\mathrm{round}}-z)\ll(N-\mathrm{shift})]$
  $\mathbf{x}^{\texttt{INT}},s_{\mathrm{x}}\leftarrow\text{act\_quantizer}(\mathbf{x})$
  $\mathbf{y}=s_{\mathrm{x}}\mathbf{W}_{\mathrm{merge}}\mathbf{x}^{\texttt{INT}}$
Appendix B Distribution Visualization of $\sigma_{\mathbf{R}}$
In Sec. 4.2, we have theoretically pointed out that there is a choice dilemma for $\sigma_{\mathbf{R}}$. Here we elaborate on its effect through distribution visualization. Specifically, we remove the VMC and use a scaling scalar to generate a too-large or too-small auxiliary variance, followed by the log2 quantization on the adaptation term. The results are shown in Fig. 8. On the one hand, setting $\sigma_{\mathbf{R}}$ too large can lead to a low correlation $\rho(\mathbf{W},\mathbf{W-R})$, which makes it hard to reconstruct $\mathbf{W}$ from $\mathbf{W-R}$ using estimator $\mathbf{W}\approx\mathcal{Q}(\mathbf{W-R})+\mathbf{R}$. On the other hand, a too small $\sigma_{\mathbf{R}}$ prevents the expectation of adaptation term converging to zero, causing few log bins to be used. In experiments, we find that the training of both settings fails to converge. By contrast, the proposed VMC can precisely control $\sigma_{\mathbf{R}}$ to allow most values of the adaptation term to be zero-neighbored, facilitating more challenging log2 quantization. Moreover, it should be noted that the too-small $\sigma_{\mathbf{R}}$ can also be regarded as an approximation of direct quantization on the zero-initialized $\mathbf{AB}$, and thus the experimental results also justify the AQS for zero-initialized $\mathbf{AB}$.
Appendix C More Implementation Details
For the subject-driven generation, we use the AdamW optimizer with a weight decay of 1e-2 and fine-tune the query, key, value, and output projection layer. The learning rate is set to 6e-5. The batch size is set to 1, and the number of training steps is 400. The rank of the LoRA is set to 4. We adopt the prior preservation strategy as Dreambooth (Ruiz et al., 2023) to generate 200 class images. For the controllable generation, we fine-tune the model for 11 epochs for Canny-to-Image tasks and 20 epochs for Landmark-to-Face and Segmentation-to-Image tasks. The learning rate is set to 1e-5 using the AdamW optimizer. The LoRA rank is set to 4. The batch size is set to 8 and the image resolution is $512\times 512$ for all three tasks. For the style-customized generation, we fine-tune the pre-trained model using the AdamW optimizer with a learning rate of 5e-5. Since it involves a larger SDXL, we chose a relatively large LoRA rank of 64 for all compared methods, since there is only one style image as well as the larger pre-trained parameters. We fine-tune for 500 steps with batch size 1. Similar to StyleDrop (Sohn et al., 2023), we only use one image as the style image and find it works well. The style images and text prompts for evaluation are given in Appendix H. The variance ratio in the variance matching control is surrogated as the value range ratio, i.e., $r=\frac{\max\{|\max(\mathbf{W})|,|\min(\mathbf{W})|\}}{\min\{|\max(\mathbf{R})|,|\min(\mathbf{R})|\}}$. We append the trainable low-rank parameters on the Query, Key, Value, and Out projection, in the attention layers (Vaswani et al., 2017) and keep all other layers frozen and quantized. The rank of LoRA is set to 4 for subject-driven generation and controllable generation, and 64 for style-customized generation.
@@ Appendix C More Implementation Details
Appendix D Additional Ablation Experiments
@@
Ablation on the LoRA rank. The low-rank $d$ in LoRA is a trade-off between performance and efficiency. A larger rank improves the adaptation ability by training more parameters but comes with larger training and storage costs, and vice versa. Here, we give the impact of different rank setups on performance in  Fig. 9. One can see that the performance generally improves as we increase the rank, but the rate of growth varies. For instance, the increased speed from rank=4 to rank=8 increases inferior to the one from rank=2 to rank=4. Moreover, increasing the rank to 16 can generally obtain better results than its lower counterpart. In practice, considering the trade-off between performance and efficiency, we select a moderate rate rank=4.
@@
The effects of variance matching control for IntLoRAMUL. In this work, we propose the variance matching control to adjust the variance of $\mathbf{R}$, so that allows the log2-quantization of the adaptation term to obtain the IntLoRASHIFT. In other words, the VMC is primarily introduced for IntLoRASHIFT. Despite we also apply the VMC to IntLoRAMUL, given the IntLoRAMUL does not require such strict constraints on the distribution shape of the adaptation term, it is interesting to investigate the influence of variance matching control on the performance of IntLoRAMUL. To this end, we adjust the smoothing factor $\alpha$ to adjust the strength of the VMC, e.g., setting $\alpha$ to zero can lead to the removal of the VMC. The results of the IntLoRAMUL under different VMC scales are shown in  Fig. 10. As one can see, despite the VMC being initially proposed for the log2-quantization, the well-structured distribution also facilitates uniform quantization. For example, when we set the $\alpha$ approaching zero, i.e., the VMC is close to being removed, and the performance of IntLoRAMUL appears similar pattern as the IntLoRASHIFT, which suffers a significant performance drop. Moreover, the performance gains gradually converge when the $\alpha>1.5$. In short, the VMC can not only allow the log2-quantization to work but also improve the performance of the uniform quantization.
@@
Figure 12: The distribution visualization of the original weights $\mathbf{W}$, the auxiliary matrix $\mathbf{R}$, and the learned low-rank weights $\mathbf{AB}$.
Appendix E Impacts from the Auxiliary Matrix
In Eq. 2 of the proposed AQS, we introduce an additional auxiliary matrix $\mathbf{R}$ to the original pre-trained weight $\mathbf{W}_{0}$ to achieve adaptation-quantization separation. However, this extra $\mathbf{R}$ potentially introduces outliers and thus causes quantization error for $\mathbf{W}$. Here, we point out that since the proposed VMC can control the range of $\mathbf{R}$ through the variance scaling factor $r=\sigma_{\mathbf{W}}/\sigma_{\mathbf{R}}$, the introduction of $\mathbf{R}$ in the AQS is ensured not result in additional outliers. For validation, we also give the distribution visualization of the original $\mathbf{W}$ and the VMC re-scaled $\mathbf{R}$ in Fig. 12. It can be seen that the range of $\mathbf{R}$ is effectively controlled within the range of $\mathbf{W}$, thus effectively avoiding the detrimental effect of additional outliers.